# MobilityGPT: Enhanced Human Mobility Modeling with a GPT model

### v2

@@ Abstract
Generative models have shown promising results in capturing human mobility characteristics and generating synthetic trajectories. However, it remains challenging to ensure that the generated geospatial mobility data is semantically realistic, including consistent location sequences, and reflects real-world characteristics, such as constraining on geospatial limits. We reformat human mobility modeling as an autoregressive generation task to address these issues, leveraging the Generative Pre-trained Transformer (GPT) architecture. To ensure its controllable generation to alleviate the above challenges, we propose a geospatially-aware generative model, MobilityGPT. We propose a gravity-based sampling method to train a transformer for semantic sequence similarity. Then, we constrained the training process via a road connectivity matrix that provides the connectivity of sequences in trajectory generation, thereby keeping generated trajectories in geospatial limits. Lastly, we proposed to construct a preference dataset for fine-tuning MobilityGPT via Reinforcement Learning from Trajectory Feedback (RLTF) mechanism, which minimizes the travel distance between training and the synthetically generated trajectories. Experiments on real-world datasets demonstrate MobilityGPT’s superior performance over state-of-the-art methods in generating high-quality mobility trajectories that are closest to real data in terms of origin-destination similarity, trip length, travel radius, link, and gravity distributions.
1 Introduction
@@
The widespread integration of location-based services and smart GPS devices, such as smartphones and watches, has made continuous monitoring of human mobility both desirable and feasible. These technologies capture diverse and detailed human movement information, with mobility trajectories representing the finest granularity of individual-level mobility characteristics. Such trajectories are crucial in various applications, including mobility modeling, commercial business analysis, and disease spread control (Siła-Nowicka et al., 2016; Changruenngam et al., 2020).
@@
Despite increasing demand for human mobility trajectory datasets, numerous challenges hinder their access and distribution (Kong et al., 2023). First, these datasets are typically collected by private companies or government agencies, hence presenting privacy concerns due to the potential disclosure of an individual’s sensitive lifestyle patterns (e.g., home or office addresses and points of interest). Legal implications, such as those outlined in the California Data Privacy Act and GDPR, highlight the importance of carefully handling spatiotemporal traces. Second, datasets owned by companies may expose proprietary business models and are often inaccessible for research purposes (Wang et al., 2019). Lastly, publicly available datasets often lack diversity or quality, with gaps in data points and intrinsic noise that significantly reduce their utility. These limitations impede the progress of urban planning or transportation research (Luca et al., 2021). Therefore, it is imperative to establish alternative trajectory data sources that are both high-quality and accessible for research purposes.
@@
Current generative models designed for human mobility exhibit multiple drawbacks: they fail to capture the sequential mobility characteristics, often lack the continuity of generated trajectories, and do not consistently adhere to geospatial constraints. Approaches based on GANs (Goodfellow et al., 2020) and VAEs (Kingma & Welling, 2013) structure data in a tabular format, which, unfortunately, fails to preserve the inherent correlations among locations. Models like LSTMs, and sequential GANs (Yu et al., 2017) struggle to maintain the smooth continuity of locations and capture realistic human mobility characteristics. A further limitation of these methods lies in the elevated mismatches observed in geospatial location sequences. In short, a successful generative model for trajectory datasets should precisely grasp the intrinsic spatial-temporal behaviors.
@@
Large language models (LLM), such as GPT (Generative Pre-trained Transformer), offer a robust foundation for sequence generation tasks and can be refined through fine-tuning to cater to distinct objectives. Trajectories and sentences share several similarities that make GPTs a promising approach for trajectory generation. Both consist of ordered sets of elements chosen from finite pools (road links and words, respectively). They exhibit semantic or spatio-temporal relationships, adhering to distinct rule systems such as language rules for sentences and geographical constraints for trajectories. Due to these parallels, the techniques developed for natural language processing can be adapted to model and generate realistic trajectories. (Musleh et al., 2022) Besides, the gravity model holds significance in human mobility modeling as it provides a structured framework for estimating and understanding the flow of interactions between different locations (Zipf, 1946). Using gravity as part of human mobility modeling could bring further advantages to synthetic trajectory generation tasks.
@@
In this study, we leverage the power of GPT for synthetic trajectory generation. However, directly training GPT on mobility data without geospatial insights would generate unrealistic sequences. To generate more realistic synthetic trajectories, we introduced several innovative methods, including (1) gravity-aware sampling, which incorporates gravity modeling of trajectory data into training updates, and (2) the use of road connectivity matrix (RCM) masking to eliminate disconnected location sequences from logits (see Figure 1). Pretraining GPT models on trajectory data captures the general sequence characteristics in an unstructured manner. We, furthermore, proposed an automated fine-tuning pipeline that improves the trajectory quality by leveraging transportation-specific metrics to evaluate and optimize the generated sequences without relying on human labeling. Our multi-objective framework, MobilityGPT, not only extends the application of GPT models to human mobility modeling but also collectively enhances the model’s capacity to capture intricate patterns and adhere to realistic geospatial constraints (e.g., trajectories consist of road segments that are indeed connected). Our contributions can be summarized as follows:
  • We propose an LLM-inspired MobilityGPT human mobility modeling approach incorporating multiple unique features with a geospatially aware pretraining method and automated fine-tuning approach without human feedback.
  • While gravity-aware sampling trains a generative model with respect to Origin-Destination (OD) pair-gravity values, the next sequence prediction is conditioned on the RCM that preserves the continuity of the generated trajectories.
  • We propose a novel method for constructing a preference dataset to fine-tune MobilityGPT using reinforcement learning, enhancing the similarity of generated trajectories in terms of travel length.
  • Through comprehensive experimental analysis using real datasets, we demonstrated that the proposed methods can generate trajectories with high fidelity while preserving essential statistical and semantic properties of human mobility.
2 Related Work
@@
Generating synthetic mobility trajectories from real data is a promising approach for protecting sensitive information (Kapp et al., 2023). Traditional model-based methods, such as those parameterizing human mobility, have limitations due to the complexity of mobility characteristics. Early data-driven approaches, employing techniques like gravity models and decision trees (Pappalardo et al., 2016; Jiang et al., 2016), face challenges in capturing sequential transitions between locations. In contrast, model-free techniques, particularly generative machine learning models, have gained traction for their ability to learn from data without external input parameters (Baratchi et al., 2014). Such generative models have been explored for human mobility modeling using various data representations, including grids in tabular format (Ouyang et al., 2018), image-like trajectory modeling (Cao & Li, 2021), and sequential grid format (Feng et al., 2020). However, these models exhibit limitations in generating trajectories that accurately capture geospatial complexity.
One promising generative model direction relies on GPS sampling. Due to mismatches, the LSTM-based trajectory generation model cannot provide continuous motion of human mobility in diverse geospatial areas (Rao et al., 2020). A similar continuous GPS trajectory generation method with a U-Net neural network using Diffusion models is proposed in (Zhu et al., 2024). Another generative direction with ML for mobility trajectories is employing two-stage training methods (Wang et al., 2021; Jiang et al., 2023; Long et al., 2023). A recent GAN-based model, relying on road links with the $A^{*}$ path-finding algorithm between regions employing two-stage GANs, has shown promise in identifying diverse paths for specified origin-destination pairs (Jiang et al., 2023). However, it falls short in generating synthetic trajectories, as the generation process initiates with an origin-destination pair from testing trajectories. Finally, sequential GAN methods have also been employed for mobility modeling with different variants (Yu et al., 2017; Feng et al., 2020). Unlike previous methods, MobilityGPT leverages the auto-regressive GPT model for trajectory generation, enabling a comprehensive exploration of spatial and temporal distributions with high-quality outputs by incorporating road links into sequence modeling.
@@
The transformer-based GPT concept has been applied to diverse applications, from vision (Chen et al., 2020), music (Banar & Colton, 2022) to network data (Kholgh & Kostakos, 2023). In addition, researchers have made great progress on various fine-tuning models for enhancing the GPT model capabilities (Zhang et al., 2019; Brown et al., 2020; Lambert et al., 2022). However, these fine-tuning models mainly rely on human feedback that evaluates the quality of the generated sequences. This study proposes a geospatially-aware GPT model, MobilityGPT, with an innovative fine-tuning approach for generating synthetic mobility trajectories.
3 Preliminaries
3.1 Discretization with Map-Matching
@@
Let $D(V,E)$ represent the road network as a weighted digraph, where the set of nodes $V$ correspond to road intersection, set of edges $E$ to roads, and weights representing link metrics, such as length of the link or traffic volume. There are two sets of trajectories: GPS trajectories and link trajectories.
@@
The map-matching process is crucial in refining GPS coordinates, which provide an approximate device location but may not consistently represent the exact travel path due to various errors. Map-matching generates an ordered set of road network links describing the user’s trajectory after accounting for the underlying road network $D(V,E)$ and GPS points (Quddus et al., 2007). Addressing location sequences in the continuous domain is challenging in conventional human mobility modeling, often resorting to a discretized grid domain. However, this grid-like representation does not entirely eliminate geospatial mismatching risks. To overcome this, our MobilityGPT model discretizes locations into road segments with map-matching, resulting in more meaningful and realistic generated trajectories.
3.2 Problem Definition
In an urban setting, a generative model is tasked with learning the intrinsic mobility characteristics from a provided set of link trajectories, $\Phi\in\Lambda=\{\Phi_{1},\Phi_{2},...,\Phi_{s}\}$. The objective is to generate synthetic trajectories that closely resemble the training trajectories in terms of various utility metrics, thereby capturing and reproducing the key features of mobility present in the input data.
4 Mobility Modeling with Generative Transformers
@@
4.1 Generative Transformers
Mobility trajectories refer to the sequential records of locations or movements over time, capturing the spatiotemporal patterns of an entity’s mobility. Their characteristics play a crucial role in human mobility modeling, providing insights into individual or collective movement behaviors.
Recent breakthroughs in NLP have proven that transformer architectures are remarkably effective in processing word sequences (Zhao et al., 2023). We have found that trajectory sequences, representing the spatiotemporal movements of individuals or objects, share similarities with sentences across four key dimensions: sequential dependencies, spatial relationships, contextual embeddings, and variable-length sequences. First, trajectory sequences exhibit sequential dependencies similar to sentences, where the order of locations matters. Additionally, trajectories involve spatial relationships between locations, just as words in a sentence convey semantic relationships. Furthermore, trajectory sequences, like sentences, can have variable lengths. Finally, trajectories benefit from contextual embeddings that consider the entire sequence, similar to the contextual understanding of words in sentences.
Given these strong similarities between trajectories and word sequences, MobilityGPT benefits two key elements of transformer architecture: self-attention and autoregressive generation. It allows efficient handling of sequential trajectory data by capturing long-range dependencies and generating contextually coherent sequences.

### v3

@@ Abstract
Generative models have shown promising results in capturing human mobility characteristics and generating synthetic trajectories. However, it remains challenging to ensure that the generated geospatial mobility data is semantically realistic, including consistent location sequences, and reflects real-world characteristics, such as constraining on geospatial limits. We reformat human mobility modeling as an autoregressive generation task to address these issues, leveraging the Generative Pre-trained Transformer (GPT) architecture. To ensure its controllable generation to alleviate the above challenges, we propose a geospatially-aware generative model, MobilityGPT. We propose a gravity-based sampling method to train a transformer for semantic sequence similarity. Then, we constrained the training process via a road connectivity matrix that provides the connectivity of sequences in trajectory generation, thereby keeping generated trajectories in geospatial limits. Lastly, we proposed to construct a preference dataset for fine-tuning MobilityGPT via Reinforcement Learning from Trajectory Feedback (RLTF) mechanism, which minimizes the travel distance between training and the synthetically generated trajectories. Experiments on real-world datasets demonstrate MobilityGPT’s superior performance over state-of-the-art methods in generating high-quality mobility trajectories that are closest to real data in terms of origin-destination similarity, trip length, travel radius, link, and gravity distributions. We release the source code and reference links to datasets at https://github.com/ammarhydr/MobilityGPT.
I Introduction
@@
The widespread integration of location-based services and smart GPS devices, such as smartphones and watches, has made continuous monitoring of human mobility both desirable and feasible. These technologies capture diverse and detailed human movement information, with mobility trajectories representing the finest granularity of individual-level mobility characteristics. Such trajectories are crucial in various applications, including mobility modeling, commercial business analysis, and disease spread control [uppoor2013generation, wang2021survey, zhou2024unified]. Understanding and modeling these mobility patterns has become increasingly critical as population growth and complex travel behaviors reshape modern transportation systems. The impacts of human mobility extend beyond transportation, influencing urban planning, environmental quality, public health, and energy consumption. Recent global events, particularly the COVID-19 pandemic, have highlighted the dynamic nature of mobility patterns and their profound societal implications [zhang2021effect, hu2021human]. This has intensified the need for sophisticated mobility modeling approaches that can capture, analyze, and predict human movement patterns at various spatial and temporal scales.
@@
Despite increasing demand for human mobility trajectory datasets, numerous challenges hinder their access and distribution [kong2023mobility, song2017multi]. First, these datasets are typically collected by private companies or government agencies, hence presenting privacy concerns due to the potential disclosure of an individual’s sensitive lifestyle patterns (e.g., home or office addresses and points of interest). Legal implications, such as those outlined in the California Consumer Privacy Act [bonta2022california] and General Data Protection Regulation [GDPR2016], highlight the importance of carefully handling spatiotemporal traces. Second, datasets owned by companies may expose proprietary business models and are often inaccessible for research purposes [wang2019urban]. Lastly, publicly available datasets often lack diversity or quality, with gaps in data points and intrinsic noise that significantly reduce their utility. These limitations impede the progress of urban planning or transportation research [luca2021survey]. Therefore, it is imperative to establish alternative trajectory data sources that are both high-quality and accessible for research purposes.
@@
Current generative models designed for human mobility exhibit multiple drawbacks: they fail to capture sequential mobility characteristics, often lack the continuity of generated trajectories, and do not consistently adhere to geospatial constraints. Approaches based on the Generative Adversarial Network (GAN) [goodfellow2020generative] and the Variational Autoencoder (VAE) [kingma2013auto] structure data in a tabular format, which unfortunately fails to preserve the inherent correlations between locations. Models like Long-Short-Term Memory (LSTM) and sequential GAN [yu2017seqgan] struggle to maintain smooth continuity of locations and capture realistic human mobility characteristics. A further limitation of these methods lies in the elevated mismatches observed in geospatial location sequences. In short, a successful generative model for trajectory datasets should precisely grasp the intrinsic spatial-temporal behaviors.
GPT (Generative Pre-trained Transformer) is a language model architecture that learns to generate coherent sequences by predicting the next token in a sequence, based on the context of previous tokens. It is trained in an autoregressive manner and excels at modeling sequential dependencies. This approach offers a robust foundation for sequence generation tasks and can be refined through fine-tuning to cater to distinct objectives.
@@
Trajectories and sentences share several similarities that make GPTs a promising approach for trajectory generation. Both consist of ordered sets of elements chosen from finite pools (road links and words, respectively). They exhibit semantic or spatiotemporal relationships, adhering to distinct rule systems such as language rules for sentences and geographical constraints for trajectories. Due to these parallels, the techniques developed for natural language processing can be adapted to model and generate realistic trajectories [musleh2022let]. In this work, we adapt the GPT architecture to the domain of human mobility by treating trajectories as sequences of spatial tokens (e.g., road links). We found that the self-attention mechanism captures intricate dependencies in human mobility sequences. This positions such models as powerful tools for overcoming challenges in human trajectory generation. Besides, the gravity model holds significance in human mobility modeling as it provides a structured framework for estimating and understanding the flow of interactions between different locations [zipf1946p]. Using gravity as part of human mobility modeling could bring further advantages to synthetic trajectory generation tasks.
@@
We leverage the power of GPT architecture for synthetic trajectory generation. However, directly training a GPT on mobility data without geospatial insights would generate unrealistic sequences. To generate more realistic synthetic trajectories, we introduced several innovative methods, including (1) gravity-aware sampling, which incorporates gravity modeling of trajectory data into training updates, and (2) the use of road connectivity matrix (RCM) masking to eliminate disconnected location sequences from logits (see Figure 1). Pretraining GPT models on trajectory data captures the general sequence characteristics in an unstructured manner. We, furthermore, proposed an automated fine-tuning pipeline that improves the trajectory quality by leveraging transportation-specific metrics to evaluate and optimize the generated sequences without relying on human labeling. Our multi-objective framework, MobilityGPT, not only extends the application of GPT models to human mobility modeling but also collectively enhances the model’s capacity to capture intricate patterns and adhere to realistic geospatial constraints (e.g., trajectories consist of road segments that are indeed connected). MobilityGPT can generate contextually rich and semantically accurate trajectories that reflect realistic human mobility without requiring prohibitive computational complexity.
Contributions: In summary, this paper addresses concerns like trajectory continuity, adherence to geospatial constraints, and computational complexity in existing models. The core of our contribution is MobilityGPT, a human mobility modeling framework that features geospatially aware pretraining and automated fine-tuning without requiring human feedback. We enhance trajectory generation by implementing gravity-aware sampling that trains our generative model using Origin-Destination (OD) pair-gravity values, while ensuring trajectory continuity through next sequence prediction conditioned on the road network through RCM. To improve trajectory quality, we developed a strategy to construct preference datasets for fine-tuning MobilityGPT via reinforcement learning, which significantly enhances travel length similarity in generated paths. Our comprehensive experimental analysis with real-world datasets demonstrates that these methods produce high-fidelity trajectories that maintain both the statistical properties and semantic characteristics essential to authentic human mobility patterns.
The remainder of the paper is organized as follows. We present related work in Section II, establish our preliminary definitions in Section III, and introduce MobilityGPT components in Section IV. After empirical evaluation in Section V, we discuss limitations in Section LABEL:s:discuss. Finally, Section LABEL:s:conclusion concludes the paper.
II Related Work
Human mobility simulation presents a significant challenge in mobility data mining due to complex geospatial and temporal correlations in movement sequences. Existing methods aim to generate realistic trajectories that reflect human behavior. Many focus on movement prediction [ying2011semantic, Qingyue2025Universal] or trajectory reconstruction [jiang2023continuous]. However, they often miss dynamic mobility patterns. They also struggle to generalize across large-scale datasets, limiting their use in comprehensive mobility modeling.
The methods for synthetic mobility trajectory generation can be broadly categorized into knowledge-driven model-based and data-driven model-free approaches. The latter predominantly leveraging machine learning techniques. Model-based methods rely on theoretical frameworks and domain-specific hypotheses, such as gravity models and spatial interaction theory, to simulate mobility patterns. Examples include microscopic traffic simulators like SUMO [lopez2018microscopic] and VISSIM [fellendorf2010microscopic]. These methods, however, often struggle to capture the complexity and variability of real-world human mobility, particularly at finer granularities.
There exists another important distinction in the trajectory literature between point-based and point-of-interest-based generative models. Point-based location sequences are dense and contain richer human movement information at high granularity. Recently, TrajGPT [hsu2024trajgpt] proposed a visitation-based sequence generation model that generates sparse location sequences based on visited locations. However, this method falls short in dense sequence prediction tasks. MobilityGPT takes a fundamentally different approach from these works, addressing the limitations of sparse sequence models by focusing on continuous, high-resolution trajectory generation that captures the complete dynamics of human movement patterns.
@@ II Related Work
Early data-driven model-free approaches, employing techniques such as gravity models, decision trees, and Markovian methods such as Mobility Markov Chains (MMC) [pappalardo2016human, jiang2016timegeo, gambs2012next], face challenges in capturing sequential transitions between locations. In contrast, generative machine learning models have gained traction for their ability to learn from data without external input parameters [baratchi2014hierarchical]. Such generative models have been explored for human mobility modeling using various data representations, including grids in tabular format [ouyang2018non], image-like trajectory modeling [cao2021generating], and sequential grid format [feng2020learning]. However, these models exhibit limitations in generating trajectories that accurately capture geospatial complexity.
One promising generative model direction relies on GPS sampling. Due to mismatches, LSTM-based trajectory generation models cannot provide continuous motion of human mobility in diverse geospatial areas [rao2020lstm]. A similar continuous GPS trajectory generation method with a U-Net neural network using Diffusion models is proposed in [zhu2024difftraj]. Sampling GPS locations without geospatial constraints diminishes the quality of synthetic trajectories. They rarely incorporate geospatial constraints and road network topologies into their training pipeline, resulting in physically impossible trajectories that ignore natural barriers and transportation infrastructure limitations when applied to new environments.
Another generative direction in ML for mobility trajectories employs two-stage training methods [wang2021large, jiang2023continuous, long2023practical]. A recent GAN-based model that relies on road links with the $A^{*}$ path-finding algorithm between regions using two-stage GANs has shown promise in identifying diverse paths for specified OD pairs [jiang2023continuous]. It falls short in generating truly synthetic trajectories, as the generation process requires initiating with an OD pair from testing trajectories. Another two-stage approach generates mobility trajectories using dual VAEs conditioned on user-level mobility characteristics learned in the second VAE model [long2023practical]. This approach is limited by its requirement for learning user-driver behaviors across multiple trips, necessitating substantial data per individual. In contrast, our MobilityGPT model learns generalist human mobility patterns without requiring personalization, making it more adaptable and efficient with available data.
Finally, sequential GAN methods have also been employed for mobility modeling with different variants [yu2017seqgan, feng2020learning]. These earlier models generate data in sequences but struggle to capture geospatial complexity and add significant computational complexity, as discussed later in our ablation studies LABEL:tab:comp_demand. Unlike previous methods, MobilityGPT leverages the autoregressive GPT model for trajectory generation. This enables a comprehensive exploration of spatial and temporal distributions with high-quality outputs by incorporating road links into sequence modeling.
@@
The GPT models, developed by [radford2018improving], transformed the field of Natural Language Processing (NLP) with the transformer’s capabilities. GPT, with its pre-training on diverse language data, demonstrated remarkable proficiency in understanding and generating coherent sequences to support various NLP tasks. The transformer-based GPT concept has been applied to diverse applications, from vision [chen2020generative], music [banar2022systematic] to network data [kholgh2023pac]. In addition, researchers have made great progress on various fine-tuning models for enhancing the GPT model capabilities [zhang2019ernie, brown2020language, lambert2022illustrating]. However, these fine-tuning models mainly rely on human feedback that evaluates the quality of the generated sequences. This study proposes a geospatially-aware GPT model, MobilityGPT, with an innovative fine-tuning approach for generating synthetic mobility trajectories.
III Preliminaries
III-A Discretization with Map-Matching
@@
Let $D(V,E)$ represent the road network as a weighted digraph, where the set of nodes $V$ corresponds to the intersection of the road, the set of edges $E$ to the roads and the weights representing the metrics of the link, such as the length of the link or traffic volume. There are two sets of trajectories: GPS trajectories and link trajectories.
@@
3) Map-matching: GPS coordinates provide an approximation of a device’s location by utilizing satellite broadcast information. Nonetheless, these coordinates may not consistently reflect the precise travel path, as they are susceptible to intrinsic and environmental errors, including satellite geometry, signal blockage, or tree cover [berber2012comparison]. Consequently, the recorded GPS locations might not align with a specific link on the road network. To address this, map-matching is employed to generate an ordered set of road network links that accurately describes the user’s trajectory, taking into account the road network $D(V,E)$ and GPS points [quddus2007current].
Addressing location sequences in the continuous domain poses challenges. Consequently, conventional human mobility modeling methods often resort to the discretized grid domain. However, relying on a grid-like discrete representation does not eliminate the risk of geospatial mismatching. For instance, random sampling from a sparse road network grid may lead to nonsensical locations in generated trajectories. In response to this limitation, our MobilityGPT model adopts a strategy of discretizing locations into road segments, enhancing the quality of generated trajectories.
III-B Problem Definition
In an urban setting, a generative model is tasked with learning the intrinsic mobility characteristics from a provided set of link trajectories, $\Phi\in\Lambda=\{\Phi_{1},\Phi_{2},...,\Phi_{s}\}$. The objective is to model human mobility using a sequential learning approach and generate synthetic trajectories that resemble training data across key utility metrics.
The sequential nature of mobility trajectories encapsulates human routing behavior, where the successive road segments influence subsequent link choice. The generative model aims to learn the conditional probability distribution $P(\phi_{k+1}|\phi_{k},\phi_{k-1},...,\phi_{k-m})$, capturing the sequential dependencies that characterize human mobility patterns. The generation process involves producing synthetic trajectories that maintain both network connectivity constraints and realistic routing behaviors observed in the training data.
IV Mobility Modeling with Generative Transformers
@@
IV-A Generative Transformers
Mobility trajectories are the sequential records of locations or movements over time, capturing the spatiotemporal patterns of an entity’s mobility. Their characteristics are crucial in human mobility modeling, providing insights into individual or collective movement behaviors.
Recent breakthroughs in NLP have proven that transformer architectures are remarkably effective in processing word sequences [zhao2023survey]. Trajectory sequences, representing the spatiotemporal movements of individuals or objects, share similarities with sentences across four key dimensions: sequential dependencies, spatial relationships, contextual embeddings, and variable-length sequences. First, trajectory sequences exhibit sequential dependencies similar to sentences, where the order of locations matters. Additionally, trajectories involve spatial relationships between locations, just as words in a sentence convey semantic relationships. Furthermore, trajectory sequences, like sentences, can have variable lengths. Finally, trajectories benefit from contextual embeddings that consider the entire sequence, similar to the contextual understanding of words in sentences.
Given these strong similarities with trajectories and word sequences, two key elements of transformers, the self-attention mechanism, and autoregressive generation, empower MobilityGPT to handle sequential data efficiently. The self-attention mechanism captures the long-range dependencies and weighs the importance of each location in the context of the entire trajectory, enabling the modeling of sequential dependencies. Besides, autoregressive sequence generation involves predicting one element at a time based on the context of preceding elements, contributing to the model’s ability to generate coherent and contextually relevant trajectories.